|
Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
FXE
SPB
HED
SQS
SXP
Sample Environment
PSPO
XO
TS
Migrated2Zulip
Archived2GitlabPages
Archived2PDF
MigratedByGroup
|
|
SCS
Photon Commissioning Team
XRO
DOC
WP74
MID
XO_mig
XO_Operation
SQS Migrated2Zulip
DET
|
|
Technical integration and coordination
|
| Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
|
Message ID: 409
Entry time: 05 Apr 2023, 09:47
|
| Author: |
Frederik Wolff-Fabris |
| Location: |
General |
| Severity: |
Information |
| Subject: |
Notes XO Integration Meeting 05.04.2022 |
|
|
General Announcement:
- Last week the template for the summer maintenance was send to the groups. In case of questions please contact M. Doering.
- UOM will not happen this Friday due to the holidays.
CTRL:
- HED setpoint issue is solved.
DA:
- Nothing to report.
EEE:
- Checking a ticket on issues with HED train-IDs; being addressed by the DOC.
VAC:
- XTD10 will be open next Tuesday 11.04 to exchange a PES turbo pump. It is a chance for additional accesses until after lunch.
- VAC did not notice effects of the power glitches.
XPD:
- Nothing to report.
DET:
- E-Pix at MID: the detectors are back online but reason for the issue is not yet understood.
ITDM:
- GPFS is up and running after the downtime yesterday (problem with one Maxwell node that was not accessing the system).
XRO:
- Not present.
LAS:
- Nothing to report.
UNSYS:
- ZZ access yesterday to replace two broken motors in SA3. Unsys still has 4 available spare motors.
- On April 11th Pb shield will be installed in SA3. All PS behind two color chicane in SA3 were already affected/exchanged and some for more than once.
SPB:
- Nothing to report.
FXE:
- Nothing to report.
MID:
- Nothing to report.
HED:
- For beam time starting on Friday the user will bring their own time-pix detector. User arrives only tomorrow afternoon and a test will be done prior Friday.
HED would like to get in touch with EEE and CTRL for support. Beamtime will start with AGIPD detector.
- Steffen H. as DRC mentions the DOC is around but may not be able to address some eventual issues over the holidays if experts would be needed; Nerea J. says that Frank Babies and Bruno Fernandes will get
in touch with Motoaki N.
SCS:
- Nothing to report.
SQS:
- Nothing to report.
SXP:
- Commissioning with beam is progressing good. |